OpenStudy (anonymous):

choose the equation of the line with a line passing through (3,-5) and perpendicular to x-2y=2

OpenStudy (anonymous):

ok so perpendicular to x-2y = 2 We have - x-2 = 2y i.e. x/2 -1 = y Now comparing with y = mx+c, m1 = 1/2 Now m1m2 = -1 for perpendicular lines. Hence slope of perp. line passing through (3,-5) m2 m2 = -1/1/2 => m2 = -2 Now if x1 = 3, y1 = -5, m2 = -2, Equation = y-y1 = m2(x-x1) => y+5 = -2(x-3) => y+5 = 6-2x i.e. 2x+y-1 = 0 is the answer.

OpenStudy (anonymous):

ok so the equation is y=1/2x-7/2 or y=-1/2x-7/2 or y=-2x+1 or y=-2x=11